OpenStudy (anonymous):

How Do We Turn 1 over 4 into a percentage?

OpenStudy (anonymous):

1/4=25/100=.25=25%

OpenStudy (anonymous):

Basically ask yourself what can i multiply the denominator to get 100. Multiply that number to the numerator (top half), and then convert the x/100 into a decimal. Move the decimal over 2 places to the left to make it a fraction.

OpenStudy (anonymous):

blah, into a percent**

OpenStudy (anonymous):

Oh I See. Thank You.